A nut mixture contains 15 lb of peanuts and 9 lb of cashews. Write the ratio in lowest terms of cashews to peanuts
Answer by TimothyLamb(4379)

c/p = 9/15
c/p = 3/5
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---